Determine whether the given quadratic function has a minimum value or maximum value. Then find the coordinates of the minimum or maximum point.
f(x)=x^2+2x-9

Answer by stanbon(75887)

Determine whether the given quadratic function has a minimum value or maximum value. Then find the coordinates of the minimum or maximum point.
f(x)=x^2+2x-9
----
Since the coefficient of x^2 is positive the function opens up, so it
has a minimum point.
----
Minimum occurs when x = -b/(2a) = -2/2 = -1
----
f(-1) = (-1)^2 + 2*-1 - 9 = 1 +2-1 = 2
---
Minimum point (-1,2)
=======================
Cheers,
Stan H.
=======================